# Graph-Informed Adversarial Modeling: Infimal Subadditivity of Interpolative Divergences

### 20 Mar 2026 v1

@@ 5 Graph-informed GANs
By contrast, interpolative divergences presented in Section 3.4–which have demonstrated strong empirical performance in generative modeling–have not been systematically studied from a subadditivity perspective. One reason is that these divergences are typically defined through primal–dual (variational) formulations (see equation 7 and equation 6), rather than through the classical representation equation 3. Moreover, the absence of a triangle inequality from divergences does not help in establishing the subadditivity bound, see e.g., Theorem 6 and its proof in Ding et al. (2021). These considerations introduce technical challenges that require different handling. Before introducing our new infimal notion, we record a simple consequence of existing subadditivity results together with the primal representation of $(f,\Gamma)$-divergences. This provides a useful benchmark but also illustrates that the resulting bounds are not yet in the form required for our later graph-local variational analysis.
Theorem 5.2.
If $(\mathcal{X},d)$ is a finite metric space, the $(f_{\mathrm{KL}},\mathrm{Lip}_{b}^{L}(\mathcal{X}))$-divergence, for any $L>0$, satisfies $\left(\frac{2\,\mathrm{diam}(\mathcal{X})}{d_{\min}}\right)$-linear subadditivity on Bayesian networks, where $\mathrm{diam}(\mathcal{X})=\max_{\bm{x},\bm{y}\in\mathcal{X}}d(\bm{x},\bm{y})$ and $d_{\min}=\min_{\begin{subarray}{c}\bm{x},\bm{y}\in\mathcal{X}\\ \bm{x}\neq\bm{y}\end{subarray}}d(\bm{x},\bm{y}).$
Proof.
By using the primal formulation of $(f,\Gamma)$-divergences given in equation 7, the subadditivity of KL-divergence and 1-Wasserstein given in Theorem 3 and Corollary 7 in Ding et al. (2021) and the fact that
$\begin{split}\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P)=\inf_{R\in\mathcal{P}(\mathcal{X})}\left\{LW_{1}(Q,R)+\mathfrak{D}_{f_{\mathrm{KL}}}(R\|P)\right\}\leq\inf_{R\in\mathcal{P}^{G}}\left\{LW_{1}(Q,R)+\mathfrak{D}_{f_{\mathrm{KL}}}(R\|P)\right\}\end{split}$
the result follows. ∎
@@ 5.1 Infimal subadditivity for Bayesian networks over a DAG
Although Ding et al. (2021) proves subadditivity for $f$-divergences in a restricted setting where the distributions $P$ and $Q$ are sufficiently close, this framework does not directly cover interpolative divergences such as $(f,\Gamma)$-divergences. The difficulty is that interpolative divergences are defined via infimal convolution and variational representations rather than via the classical density-based form of $f$-divergences. This suggests replacing classical subadditivity by a weaker notion that is tailored to optimization over structured model classes. Motivated by this gap, we introduce $\alpha$-linear generalized infimal subadditivity and additivity on Bayesian networks, formalized in the next definition.
@@
$\inf_{P\in\mathcal{P}(\mathcal{X})}\delta(Q,P)\leq\inf_{P\in\mathcal{P}(\mathcal{X})}\alpha\sum_{i=1}^{n}\delta_{i}\left(Q_{i\,\cup\,\mathrm{Pa}(i)}\,,P_{i\,\cup\,\mathrm{Pa}(i)}\right).$ (16)
For the case $\delta=\delta_{1}=\cdots=\delta_{n}$, we say that $\delta$ satisfies infimal subadditivity on Bayesian networks. Furthermore we say that $\delta$ satisfies $\alpha$-linear generalized infimal additivity with respect to $\delta_{1},\dots,\delta_{n}$ on Bayesian networks if equation 16 holds with equality.
Although the optimization problem in equation 16 may be viewed as graph-constrained in spirit, the infima are taken over all $P\in\mathcal{P}(\mathcal{X})$. This is important for applications as even when the target $Q$ factorizes according to $G$, it is generally difficult to enforce the same structural constraint on the optimizer $P^{\ast}$ during training. The usefulness of infimal subadditivity is therefore that it replaces a difficult global optimization problem by a surrogate built from family-level discrepancies, without requiring the optimizer itself to be explicitly parameterized as an element of $\mathcal{P}^{G}$.
We now show that this principle applies to the main discrepancy classes considered in this paper. Under suitable assumptions, $(f,\Gamma)$-divergences, $\Gamma$-IPMs, and proximal optimal transport divergences all admit graph-informed upper bounds in terms of local family objectives, and in the additive regime, these bounds become exact. We begin with the interpolative case $(f,\Gamma)$, which provides the basic template for the later results.
@@
Let $G$ be an arbitrary DAG and $L,L_{1},\dots,L_{n}>0$. Suppose that $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}_{i\cup\mathrm{Pa}(i)}$ for all $i\in V$. Then, $(f,\Gamma)$-divergences satisfy $\frac{1}{n}$-linear generalized infimal subadditivity given in equation 16 with
@@
The proof of the theorem relies on a sequence of data-processing–type inequalities involving the operators $I_{i,\mathrm{Pa}(i)}$, together with the corresponding assumptions $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}_{i\,\cup\,\mathrm{Pa}(i)}$ for all $i\in V$. To clarify the main ideas and highlight the role played by these operators, we first present the proof for the simple DAG
$X_{1}\longrightarrow X_{2}\longrightarrow X_{3}\,$, (17)
while the proof for a Bayesian network defined on an arbitrary DAG comes next. Let $\gamma\in\Gamma^{L}$, we begin with computing the expectation of $I_{1,2}[\gamma]$ with respect to $Q_{\{1,2\}}$. A direct calculation gives us
$E_{Q_{\{1,2\}}}\left[I_{1,2}[\gamma]\right]=\int_{\mathcal{X}_{1}}\int_{\mathcal{X}_{2}}\int_{\mathcal{X}_{3}}\gamma(x_{1},x_{2},x_{3})Q_{\{1,2\}}(dx_{1},dx_{2})Q_{3\mid 2}(dx_{3})=E_{Q}[\gamma].$ (18)
For $P_{\{1,2\}}\in\mathcal{P}(\mathcal{X}_{\{1,2\}})$ (we remind that $\mathcal{X}_{\{1,2\}}=\mathcal{X}_{1}\times\mathcal{X}_{2}$), for any function $f$ as defined in Section 3.1 and for $\nu\in\mathbb{R}$, the convexity $f^{*}$, Jensen’s inequality, and equation 13 give:
$E_{P_{\{1,2\}}}\left[f^{*}(I_{1,2}[\gamma]-\nu)\right]\leq E_{P_{\{1,2\}}}\left[I_{1,2}[f^{*}(\gamma-\nu)]\right]=E_{P_{\{1,2\}}\otimes Q_{3\mid 2}}[f^{*}(\gamma-\nu)]\,.$ (19)
where $P_{\{1,2\}}\otimes Q_{3\mid 2}\in\mathcal{P}_{2}^{G}$ is defined in equation 12.
@@
$E_{Q_{\{1,2\}}}\left[I_{1,2}[\gamma]\right]-\inf_{\nu\in\mathbb{R}}\left\{\nu+E_{P_{\{1,2\}}}\left[f^{*}(I_{1,2}[\gamma-\nu])\right]\right\}$
$\geq E_{Q}[\gamma]-\inf_{\nu\in\mathbb{R}}\left\{\nu+E_{P_{\{1,2\}}\otimes Q_{3\mid 2}}[f^{*}(\gamma-\nu)]\right\}\,.$
@@
$\mathfrak{D}_{f}^{I_{1,2}[\Gamma^{L}]}(Q_{\{1,2\}}\|P_{\{1,2\}})\geq\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{\{1,2\}}\otimes Q_{3\mid 2})\,$, (20)
@@
$\mathfrak{D}_{f}^{I_{1}[\Gamma^{L}]}(Q_{\{1\}}\|P_{\{1\}})\geq\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{\{1\}}\otimes Q_{3\mid 2}\otimes Q_{2\mid 1})\,$, (21)
@@
$\mathfrak{D}_{f}^{I_{2,3}[\Gamma^{L}]}(Q_{\{2,3\}}\|P_{\{2,3\}})\geq\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{\{2,3\}}\otimes Q_{1\mid 2})\,.$ (22)
@@
$c=\inf_{P\in\mathcal{P}(\mathcal{X})}\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P)\,;$
since $\mathcal{P}^{G}_{i}\subset\mathcal{P}^{G}\subset\mathcal{P}(\mathcal{X})$, in particular it holds that
$\mathfrak{D}_{f}^{I_{1}[\Gamma^{L}]}(Q_{\{1\}}\|P_{\{1\}})+\mathfrak{D}_{f}^{I_{1,2}[\Gamma^{L}]}(Q_{\{1,2\}}\|P_{\{1,2\}})+\mathfrak{D}_{f}^{I_{2,3}[\Gamma^{L}]}(Q_{\{2,3\}}\|P_{\{2,3\}})\geq 3c\,.$ (23)
Then we consider infimum over $P\in\mathcal{P}^{G}$ on both sides yields equation 16. Furthermore, from equation 20–equation 23 and the fact that $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}$ for all $i\in V$, we obtain
$\mathfrak{D}_{f}^{\Gamma^{L_{1}}}(Q\|P_{\{1\}}\otimes Q_{3\mid 2}\otimes Q_{2\mid 1})+\mathfrak{D}_{f}^{\Gamma^{L_{2}}}(Q\|P_{\{1,2\}}\otimes Q_{3\mid 2})+\mathfrak{D}_{f}^{\Gamma^{L_{3}}}(Q\|P_{\{2,3\}}\otimes Q_{1\mid 2})\geq 3c\,.$ (24)
@@
$E_{P_{i\,\cup\,\mathrm{Pa}(i)}}\left[I_{i,\mathrm{Pa}(i))}[\gamma]\right]=\int_{\mathcal{X}}\gamma(\bm{x})\tilde{P}(d\bm{x})=E_{\tilde{P}}[\gamma]$
where $\tilde{P}=Q_{V^{(i)}\mid i\,\cup\,\mathrm{Pa}(i)}\otimes\,P_{i\,\cup\,\mathrm{Pa}(i)}\otimes\prod_{j=i+1}^{n}Q_{j\mid\mathrm{Pa}(j)}\in\mathcal{P}_{i}^{G}$ and
$E_{Q_{i\,\cup\,\mathrm{Pa}(i)}}[f^{*}(I_{i,\mathrm{Pa}(i))}[\gamma]-\nu)]\leq E_{Q}[f^{*}(\gamma-\nu)]\,.$
The other direction given in equation 25 remains unchanged, and combining both inequalities completes the proof. ∎
Remark 5.5.
In Theorem 5.4, the infimum can also be taken over $P\in\mathcal{P}^{G}$; this follows directly from the proof. In practice, however, enforcing the graph structure directly in the generator class is challenging. For this reason, we retain the larger space $\mathcal{P}(\mathcal{X})$.
@@
Figure 2(a) compares a graph-agnostic GAN, whose monolithic discriminator acts on the full joint distribution, with a graph-informed GAN built from family-level discriminator classes on $\{i\cup\mathrm{Pa}(i)\}_{i=1}^{8}$. We report both the classical setting, based on unrestricted $f$-divergence objectives, and the interpolative setting, based on constrained $(f,\Gamma)$-divergence objectives when $f=f_{\mathrm{KL}}$ and $\Gamma=\mathrm{Lip}_{b}$. In the interpolative setting, the graph-informed surrogate closely tracks the best achieved global objective and exhibits substantially lower variability across seeds. By contrast, in the classical setting the graph-agnostic objective is markedly more variable, reflecting the greater difficulty of the full-joint discrimination problem.
@@
Let $G$ be an arbitrary DAG and $L,L_{1},\dots,L_{n}>0$. Suppose that $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}_{i\cup\mathrm{Pa}(i)}$ for all $i\in V$. Then, IPMs satisfy $\frac{1}{n}$-linear generalized infimal subadditivity given in equation 16 with
@@
Let $c$ be any cost function on $\mathcal{X}$. Then, proximal optimal transport divergences satisfy $\frac{1}{n}$-linear generalized infimal subadditivity given in equation 16 with
@@
$c_{i}(\bm{x}_{i\,\cup\,\mathrm{Pa}(i)},\bm{x}^{\prime}_{i\,\cup\,\mathrm{Pa}(i)})=\int_{\mathcal{X}_{[i\,\cup\,\mathrm{Pa}(i)]^{c}}}\int_{\mathcal{X}_{[i\,\cup\,\mathrm{Pa}(i)]^{c}}}\!\!c(\bm{x},\bm{x}^{\prime})Q^{(i)}(d\bm{x}_{[i\,\cup\,\mathrm{Pa}(i)]^{c}})Q^{(i)}(d\bm{x}^{\prime}_{[i\,\cup\,\mathrm{Pa}(i)]^{c}})\,$,
where $Q^{(i)}=Q_{V^{(i-1)}\mid\mathrm{Pa}(i)}\otimes\prod_{j=i+1}^{n}Q_{j\mid\mathrm{Pa}(j)}$ is the marginal distribution of $\bm{X}_{[i\,\cup\,\mathrm{Pa}(i)]^{c}}$. Moreover, the proximal optimal transport divergences made by $c(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|_{p}^{p}=\sum_{i=1}^{n}\|x_{i}-y_{i}\|_{p}^{p}$ (i.e., $p$-Wasserstein distance) and KL-divergence, satisfy $\frac{1}{n}$-infimal additivity with $\delta_{i}=\mathfrak{D}_{\rm{KL},\varepsilon}^{c_{i}}$ with $c_{i}(x_{i},y_{i})=\|x_{i}-y_{i}\|_{p}^{p}$, for all $i\in V$ .

### 02 Apr 2026 v2

@@ 5 Graph-informed GANs (GiGANs)
By contrast, interpolative divergences presented in Section 3.4—which have demonstrated strong empirical performance in generative modeling—have not been systematically studied from the viewpoint of subadditivity. A key difficulty is that these discrepancies are defined through primal–dual variational formulations, such as equation 8 and equation 7, rather than through the classical density-based representation equation 3. In particular, the standard arguments used to prove subadditivity for classical divergences do not transfer directly. The lack of a triangle inequality for divergences is an example; see Theorem 6 and its proof in Ding et al. (2021).
To motivate the need for a different notion, we present two simple calculations. The first recalls the mechanism behind subadditivity for the KL divergences, as proven in (Ding et al., 2021, Supplement §8.2). The second shows why the same argument breaks down for interpolative divergences.
To illustrate the first point, consider the simple chain,
$X_{1}\longrightarrow X_{2}\longrightarrow X_{3}\,.$ (18)
For any $P,Q\in\mathcal{P}^{G}$, the key observation is that
$\mathfrak{D}_{f_{\mathrm{KL}}}(Q\|P)$ $=E_{Q}\left[\log\left(\frac{dQ}{dP}\right)\right]$ (19)
$=E_{Q_{12}}\left[\log\left(\frac{dQ_{12}}{dP_{12}}\right)\right]+E_{Q_{23}}\left[\log\left(\frac{dQ_{23}}{dP_{23}}\right)\right]-E_{Q_{2}}\left[\log\left(\frac{dQ_{2}}{dP_{2}}\right)\right]$
$\leq\mathfrak{D}_{f_{\mathrm{KL}}}(Q_{12}\|P_{12})+\mathfrak{D}_{f_{\mathrm{KL}}}(Q_{23}\|P_{23})$
$\leq\mathfrak{D}_{f_{\mathrm{KL}}}(Q_{1}\|P_{1})+\mathfrak{D}_{f_{\mathrm{KL}}}(Q_{12}\|P_{12})+\mathfrak{D}_{f_{\mathrm{KL}}}(Q_{23}\|P_{23})$
because $P=\frac{P_{12}P_{23}}{P_{2}}$ and $Q=\frac{Q_{12}Q_{23}}{Q_{2}}$. Thus, the logarithm of the likelihood ratio decomposes into local terms, with the overlap corrected by the marginal on $X_{2}$. As shown in Ding et al. (2021), this argument extends to arbitrary DAGs by induction.
The situation is different for interpolative divergences. Consider an $(f,\Gamma)$-divergence with $f=f_{\mathrm{KL}}$ and $\Gamma=\mathrm{Lip}^{1}_{b}(\mathcal{X})$, and let $R^{*}$ denote the optimal intermediate measure that solves the infimal convolution equation 8. By Remark 4.12 of Dupuis and Mao (2022),
$\mathfrak{D}_{f_{\mathrm{KL}}}^{\Gamma}(Q\|P)=E_{Q}\left[\log\left(\frac{dR^{*}}{dP}\right)\right]\,.$ (20)
At first sight, equation 20 resembles the KL identity equation 19. The crucial difference, however, is that the interpolative divergence involves the optimal intermediate measure $R^{*}$ in place of $Q$. Moreover, equation 9 gives,
$dR^{*}\propto(f_{\rm{KL}}^{*})^{\prime}(\gamma^{*})dP=e^{\gamma^{*}-1}dP\,.$
Using the factorization of $P$, we obtain
Now consider the conditional distribution of $X_{3}$ under $R^{*}$. A direct calculation gives
$R^{*}_{3\mid\mathrm{Pa}^{*}(3)}(dx_{3})=\frac{e^{\gamma^{*}(x_{1},x_{2},x_{3})-1}\,P_{3\mid 2}(dx_{3}\mid x_{2})}{\int e^{\gamma^{*}(x_{1},x_{2},z)-1}\,P_{3\mid 2}(dz\mid x_{2})}.$
This expression reveals that, in general, the conditional distribution of $X_{3}$ under $R^{*}$ depends on both $x_{1}$ and $x_{2}$. Hence,
$\mathrm{Pa}^{*}(3)\supseteq\{2\}=\mathrm{Pa}(3)\,.$
In other words, the optimal intermediate measure $R^{*}$ need not preserve the conditional independence structure of the original graph. This prevents the local likelihood terms from being paired according to the families of $G$ as in the first calculation for the classical KL divergence.
These calculations suggest that, for interpolative divergences, classical subadditivity is generally too strong a requirement. In the next subsection, we introduce a weaker notion of subadditivity, formulated at the level of the optimization problem itself, that will allow us to incorporate this structure.
@@ 5.1 Infimal subadditivity for Bayesian networks over a DAG
The calculations in the previous section indicate that classical subadditivity fails for $(f_{\mathrm{KL}},\Gamma)$-divergences. More generally, for any $f\neq f_{\mathrm{KL}}$, subadditivity continues to fail even under the closeness assumptions considered in Ding et al. (2021). The difficulty stems from the fact that interpolative divergences are defined through infimal convolution and variational representations, rather than through the classical density-based formulation of $f$-divergences. As a result, the arguments used to establish subadditivity in the classical setting no longer apply. This suggests that classical subadditivity must be replaced by a weaker notion that is compatible with optimization over structured model classes. Motivated by this gap, we introduce the notion of $\alpha$-linear generalized infimal subadditivity (and additivity) on Bayesian networks, formalized in the next definition.
@@
$\inf_{P\in\mathcal{M}}\delta(Q,P)\leq\inf_{P\in\mathcal{M}}\alpha\sum_{i=1}^{n}\delta_{i}\left(Q_{i\,\cup\,\mathrm{Pa}(i)}\,,P_{i\,\cup\,\mathrm{Pa}(i)}\right)\,$, (21)
with $\mathcal{M}\subseteq\mathcal{P}(\mathcal{X})$. For the case $\delta=\delta_{1}=\cdots=\delta_{n}$, we say that $\delta$ satisfies infimal subadditivity on Bayesian networks. Furthermore, we say that $\delta$ satisfies $\alpha$-linear generalized infimal additivity with respect to $\delta_{1},\dots,\delta_{n}$ on Bayesian networks if equation 21 holds with equality.
The main classes of probabilities that we consider for $\mathcal{M}$ are $\mathcal{P}(\mathcal{X})$, parametric families, and $\mathcal{P}^{G}$. From a purely mathematical perspective, the infimal formulation when $\mathcal{M}=\mathcal{P}(\mathcal{X})$ is trivial, as it is minimized by $P=Q$. However, this viewpoint is not relevant in learning settings, where the target distribution $Q$ is only accessible through samples and the optimization over all probability measures is intractable. Therefore, this is important for applications as even when the target $Q$ factorizes according to $G$, it is generally difficult to enforce the same structural constraint on the optimizer $P^{\ast}$ during training. In practice, the usefulness of infimal subadditivity is therefore that it replaces a difficult global optimization problem by a surrogate built from family-level discrepancies, without requiring the optimizer itself to be explicitly parameterized as an element of $\mathcal{P}^{G}$.
We now show that this principle applies to the main discrepancy classes considered in this paper. Under suitable assumptions, $(f,\Gamma)$-divergences, $\Gamma$-IPMs, and proximal optimal transport divergences all admit graph-informed upper bounds in terms of local family objectives, and in the additive regime, these bounds become exact. The next theorem provides a theoretical justification for replacing an intractable global optimization problem with a set of tractable local learning problems. We begin with the interpolative case $(f,\Gamma)$, which provides the basic template for the later results.
@@
Let $G$ be an arbitrary DAG and $L,L_{1},\dots,L_{n}>0$. Suppose that $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}_{i\cup\mathrm{Pa}(i)}$ for all $i\in V$. Then, $(f,\Gamma)$-divergences satisfy $\frac{1}{n}$-linear generalized infimal subadditivity given in equation 21 with $\mathcal{M}=\mathcal{P}(\mathcal{X})$ or $\mathcal{P}^{G}$ and
@@
The proof of the theorem relies on a sequence of data-processing–type inequalities involving the operators $I_{i,\mathrm{Pa}(i)}$, together with the corresponding assumptions $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}_{i\,\cup\,\mathrm{Pa}(i)}$ for all $i\in V$. To clarify the main ideas and highlight the role played by these operators, we first present the proof for the simple DAG equation 18, while the proof for a Bayesian network defined on an arbitrary DAG comes next. Let $\gamma\in\Gamma^{L}$, we begin with computing the expectation of $I_{1,2}[\gamma]$ with respect to $Q_{\{1,2\}}$. A direct calculation gives us
$E_{Q_{12}}\left[I_{1,2}[\gamma]\right]=\int_{\mathcal{X}_{1}}\int_{\mathcal{X}_{2}}\int_{\mathcal{X}_{3}}\gamma(x_{1},x_{2},x_{3})Q_{12}(dx_{1},dx_{2})Q_{3\mid 2}(dx_{3})=E_{Q}[\gamma].$ (22)
For $P_{12}\in\mathcal{P}(\mathcal{X}_{\{1,2\}})$ (we remind that $\mathcal{X}_{\{1,2\}}=\mathcal{X}_{1}\times\mathcal{X}_{2}$), for any function $f$ as defined in Section 3.1 and for $\nu\in\mathbb{R}$, the convexity $f^{*}$, Jensen’s inequality, and equation 15 give:
$E_{P_{12}}\left[f^{*}(I_{1,2}[\gamma]-\nu)\right]\leq E_{P_{12}}\left[I_{1,2}[f^{*}(\gamma-\nu)]\right]=E_{P_{12}\otimes Q_{3\mid 2}}[f^{*}(\gamma-\nu)]\,.$ (23)
where $P_{12}\otimes Q_{3\mid 2}\in\mathcal{P}_{2}^{G}$ is defined in equation 14.
@@
$E_{Q_{12}}\left[I_{1,2}[\gamma]\right]-\inf_{\nu\in\mathbb{R}}\left\{\nu+E_{P_{12}}\left[f^{*}(I_{1,2}[\gamma]-\nu)\right]\right\}$
$\geq E_{Q}[\gamma]-\inf_{\nu\in\mathbb{R}}\left\{\nu+E_{P_{12}\otimes Q_{3\mid 2}}[f^{*}(\gamma-\nu)]\right\}\,.$
@@
$\mathfrak{D}_{f}^{I_{1,2}[\Gamma^{L}]}(Q_{12}\|P_{12})\geq\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{12}\otimes Q_{3\mid 2})\,$, (24)
@@
$\mathfrak{D}_{f}^{I_{1}[\Gamma^{L}]}(Q_{1}\|P_{1})\geq\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{1}\otimes Q_{3\mid 2}\otimes Q_{2\mid 1})\,$, (25)
@@
$\mathfrak{D}_{f}^{I_{2,3}[\Gamma^{L}]}(Q_{23}\|P_{23})\geq\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{23}\otimes Q_{1\mid 2})\,.$ (26)
@@
$c=\inf_{P\in\mathcal{M}}\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P)\,;$
since $\mathcal{P}^{G}_{i}\subset\mathcal{M}=\mathcal{P}^{G}$ or $\mathcal{P}(\mathcal{X})$ (and $\mathcal{P}^{G}\subset\mathcal{P}(\mathcal{X})$), in particular it holds that
$\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{1}\otimes Q_{3\mid 2}\otimes Q_{2\mid 1})+\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{12}\otimes Q_{3\mid 2})+\mathfrak{D}_{f}^{\Gamma^{L}}(Q\|P_{23}\otimes Q_{1\mid 2})\geq 3c\,.$ (27)
By using the assumption that $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}$ for all $i\in V$, we obtain
$\mathfrak{D}_{f}^{\Gamma^{L_{1}}}(Q_{1}\|P_{1})+\mathfrak{D}_{f}^{\Gamma^{L_{2}}}(Q_{12}\|P_{12})+\mathfrak{D}_{f}^{\Gamma^{L_{3}}}(Q_{23}\|P_{23})\geq 3c\,.$ (28)
Then we consider infimum over $P\in\mathcal{P}(\mathcal{X})$ on both sides yields equation 21.
@@
$E_{Q_{i\,\cup\,\mathrm{Pa}(i)}}\left[I_{i,\mathrm{Pa}(i)}[\gamma]\right]=\int_{\mathcal{X}}\gamma(\bm{x})Q(d\bm{x})=E_{Q}[\gamma]$
and
$E_{P_{i\,\cup\,\mathrm{Pa}(i)}}[f^{*}(I_{i,\mathrm{Pa}(i)}[\gamma]-\nu)]\leq E_{\tilde{P}}[f^{*}(\gamma-\nu)]$
where $\tilde{P}=Q_{V^{(i)}\mid i\,\cup\,\mathrm{Pa}(i)}\otimes\,P_{i\,\cup\,\mathrm{Pa}(i)}\otimes\prod_{j=i+1}^{n}Q_{j\mid\mathrm{Pa}(j)}\in\mathcal{P}_{i}^{G}$. The other direction given in equation 29 remains unchanged, and combining both inequalities completes the proof. When the infimum is be taken over $P\in\mathcal{P}^{G}$, the proof follows the same steps as above. ∎
Remark 5.4 (Computational tractability).
In practice enforcing the graph structure $\mathcal{P}^{G}$ directly in the generator class is challenging. Therefore, the theorem allows to construct tractable surrogate objectives that approximate or control the global divergence without requiring a graphical factorization on the model $P$. For this reason, we retain the larger space $\mathcal{P}(\mathcal{X})$. The graph structure is used to define a collection of local objectives that collectively control the global divergence.
@@
Figure 2(a) compares a graph-agnostic GAN, whose monolithic discriminator acts on the full joint distribution, with a GiGAN built from family-level discriminator classes on $\{i\cup\mathrm{Pa}(i)\}_{i=1}^{8}$. We report both the classical setting, based on unrestricted $f$-divergence objectives, and the interpolative setting, based on constrained $(f,\Gamma)$-divergence objectives when $f=f_{\mathrm{KL}}$ and $\Gamma=\mathrm{Lip}_{b}$. In the interpolative setting, the graph-informed surrogate closely tracks the best achieved global objective and exhibits substantially lower variability across seeds. By contrast, in the classical setting the graph-agnostic objective is markedly more variable, reflecting the greater difficulty of the full-joint discrimination problem.
@@
Let $G$ be an arbitrary DAG and $L,L_{1},\dots,L_{n}>0$. Suppose that $I_{i,\mathrm{Pa}(i)}[\Gamma^{L}]\subset\Gamma^{L_{i}}_{i\cup\mathrm{Pa}(i)}$ for all $i\in V$. Then, IPMs satisfy $\frac{1}{n}$-linear generalized infimal subadditivity given in equation 21 with $\mathcal{M}=\mathcal{P}(\mathcal{X})$ or $\mathcal{P}^{G}$ and
@@
Let $c$ be any cost function on $\mathcal{X}$. Then, proximal optimal transport divergences satisfy $\frac{1}{n}$-linear generalized infimal subadditivity given in equation 21 with $\mathcal{M}=\mathcal{P}(\mathcal{X})$ or $\mathcal{P}^{G}$ and
@@
$c_{i}(\bm{x}_{i\,\cup\,\mathrm{Pa}(i)},\bm{x}^{\prime}_{i\,\cup\,\mathrm{Pa}(i)})$ $=$ $\int_{\mathcal{X}_{A}\times\mathcal{X}_{A}}c(\bm{x},\bm{x}^{\prime})\prod_{\bm{y}=\bm{x},\bm{x}^{\prime}}\Big(Q(d\bm{y}_{V^{(i-1)}}|\bm{y}_{\mathrm{Pa}(i)})\prod_{j=i+1}^{n}Q(dy_{j}|\bm{y}_{\mathrm{Pa}(j)})\Big)\,$,
where $A=[i\,\cup\,\mathrm{Pa}(i)]^{c}$.